OpenStudy (anonymous):

simplify: -(3a+4b)+5a-4b

DivineSolar (divinesolar):

Do you even know how to start completing this problem?

OpenStudy (anonymous):

to be completly honest no

DivineSolar (divinesolar):

Contact your teacher first please.

OpenStudy (anonymous):

i do online schooling (:

OpenStudy (anonymous):

so my only help is online to be honest

DivineSolar (divinesolar):

FLVS correct?

OpenStudy (anonymous):

2a−8b

OpenStudy (skullpatrol):

$$\Huge -a=-1a$$

OpenStudy (anonymous):

thank you blueish but i still need to learn how to solve it

OpenStudy (anonymous):

kk I'll give you steps

OpenStudy (onepieceftw):

I'd like to take a crack at it. the negative sign before the parentheses mean you multiply the expression in the parentheses by -1

OpenStudy (onepieceftw):

Then just add like terms

OpenStudy (skullpatrol):

replace a with 3a +4b

OpenStudy (anonymous):

ohh. i feel bad for only giving one of you best response because you all are so helpful

OpenStudy (anonymous):

distribute =−3a+−4b+5a+−4b combine Like Terms =−3a+−4b+5a+−4b =(−3a+5a)+(−4b+−4b) =2a+−8b answer =2a−8b

OpenStudy (anonymous):

Thats how you do it hope i helped c;

OpenStudy (skullpatrol):

$$\Huge -(3a+4)=-1(3a+4b)$$

OpenStudy (anonymous):

MK, I'm going now if you need more help pm me c;

OpenStudy (skullpatrol):

$$\Huge a(b+c)=ab+ac$$

OpenStudy (anonymous):

thanks all (:

OpenStudy (anonymous):

I do online schooling to i know how you feel... :(